OpenStudy (snivvles.com):

find solutions between 0 and 2pi 7 sin2x - 14 sin x + 2 = -5 will fan and medal

OpenStudy (snivvles.com):

so far i've subtracted 2 from each side and divided by 7 to get sin^2x-2sinx=1

OpenStudy (snivvles.com):

@enchanted_bubbles @Peaches15

Nnesha (nnesha):

divided by 7 ?? why ?

Nnesha (nnesha):

this is like quadratic equation let sinx = x

Nnesha (nnesha):

\[\huge\rm 7x^2-14x+7=0\] now find GCF and then factor it let sinx = x

OpenStudy (snivvles.com):

to simplify the equation, because all three numbers where divisible by 7

OpenStudy (snivvles.com):

oh

Nnesha (nnesha):

ohh right you mean GCF

OpenStudy (snivvles.com):

i got x^2 -2x=0

Nnesha (nnesha):

correct so you will get \[\huge\rm \frac{ 7x^2 }{ 7}-\frac{ 14x }{ 7}+\frac{ 7 }{ 7 }=0\]

Nnesha (nnesha):

7/7 = ??

OpenStudy (snivvles.com):

1

Nnesha (nnesha):

yes right so it's x^2 -2x +1 =0

OpenStudy (snivvles.com):

OH right

Nnesha (nnesha):

after factor it ot 7(x^2-2x+1)=0 now set it equal to zero \[7\cancel{=}0 ~~~~x^2-2x+1=0\] now factor x^2-2x+1=0

OpenStudy (snivvles.com):

ok gimme a sec

Nnesha (nnesha):

:=)

OpenStudy (snivvles.com):

(x+2) (x-1)?

Nnesha (nnesha):

nope 2 times -1 = -2 but it suppose to be one find two number when you multiply them you should get 1 and when you add or subtract them you should get -2

OpenStudy (snivvles.com):

(x+2)(x-1)=0

OpenStudy (snivvles.com):

ok

Nnesha (nnesha):

easy method to factor quadratic equation find two number when you multiply them you should get product of AC and when you add or subtract them you should get middle term a=leading coefficient b=middle term c=constant term

OpenStudy (snivvles.com):

hang on a sec

OpenStudy (snivvles.com):

(x-1)^2

OpenStudy (snivvles.com):

?

Nnesha (nnesha):

yep right -1 times -1 = 1 -1-1=-2

Nnesha (nnesha):

now change it back to sinx you get (sinx -1)^2=0 solve for sin x

OpenStudy (snivvles.com):

I need to find the answers in between (0,2pi)

OpenStudy (snivvles.com):

using the unit circle

Nnesha (nnesha):

yep so sinx -1= 0 solve for sinx sinx =1 right ? use unit circle to find radian where sinx = 1

OpenStudy (snivvles.com):

ok that's what i thought thank you so much!

Nnesha (nnesha):

my pleasure!